OpenStudy (anonymous):

Evaluate the function f(x) = −7x^2 + 10x when x = 2/3.

OpenStudy (anonymous):

\[-7(\frac{2}{3})^2+10\times (\frac{2}{3})\]

OpenStudy (anonymous):

i know that but im confused at the part where you square 2/3 cause when you square 2/3, doesn't it just go back to its original form?

OpenStudy (anonymous):

\[\left(\frac{2}{3}\right)^2=\frac{4}{9}\]

OpenStudy (anonymous):

Ohh. I see it now. I got it mixed up with my other problem. thank you!